OpenStudy (anonymous):

A=1/2absin(theta), How is dA/dt related to d(theta)/dt and da/dt if only b is constant? Answer: dA/dt = 1/2*ab*cos(theta)*d(theta)/dt + 1/2*b*sin(theta)*da/dt How do i solve this

OpenStudy (anonymous):

I will use x instead of theta. We have A = 1/2 a b sin x. Then dA = 1/2 ( b sin x da + a sin x db + a b cos x dx) Therefore dA/dt = 1/2 ( b sin x da/dt + a sin x db/dt + a b cos x dx/dt)

OpenStudy (anonymous):

do u understand?

OpenStudy (anonymous):

No i dont

OpenStudy (anonymous):

That the best I can explain

OpenStudy (anonymous):

so i set up my problem dA/dt=1/2 b d/dt [a*sin(theta)] is that right